OpenStudy (zaxoanl):

a 75 kg physics student rides in the elevator. find the apparent weight of the student if: a. the elevator is moving vertically up at constant velocity of 5.0 J m/s b. the elevator decelerates as it moves up at 3.5m/s^2

OpenStudy (michele_laino):

case a) since the elevator is moving with constant velocity, then it is an inertial reference system, so the weight is given by the same formula, namely: P=m*g=...?

OpenStudy (michele_laino):

case b), on our student is acting fictitious force, due to the acceleration of the elevator. Here is the situation:  so we can write this vector equation: \[{{\mathbf{W}}_{{\mathbf{APPARENT}}}} = M{\mathbf{g}} - M{\mathbf{a}}\] from which we have the subsequent scalar equation: \[\left| {{{\mathbf{W}}_{{\mathbf{APPARENT}}}}} \right| = Mg + Ma\] for the magnitude of the apparent weight